# PPFPL: Cross-silo Privacy-preserving Federated Prototype Learning Against Data Poisoning Attacks on Non-IID Data

### v2

@@ 5.2.2 Prototype Handling
5.3 Secure Aggregation
To resist local prototypes submitted by malicious clients without compromising privacy, we design a secure aggregation protocol across two servers to aggregate global prototypes. The protocol consists of normalization verification and secure two-party computation.
5.3.1 Normalization verification
Since malicious participants may bypass the normalization stage to amplify the impact of their local prototypes, the two servers need to verify that encrypted local prototypes are normalized. Specifically, Aggregator calculates the inner product $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ for each local prototype, and sends the results to Verifier, where $\cdot$ denotes the inner product. Then, the Verifier decrypts $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ using its secret key $Skv$, and checks whether the inner product $\|\widetilde{\textit{{c}}}^{k}_{m,t}\|^{2}$ equals 1. If the inner product of prototype from client $m$ is not equal to 1, which indicates that its local prototype does not normalized, client $m$ is removed from the set of clients $\mathcal{S}$. After validation, Verifier sends the set $\mathcal{S}$ to Aggregator.
5.3.2 Secure Two-party Computation
@@ 19       Distribute global prototype $\llbracket\textit{{C}}^{k}_{t+1}\rrbracket_{Pkx}$ to clients;
where $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ denotes the trusted prototype for class $k$ in $(t+1)$-th communication round. $\textit{{C}}^{\prime k}_{t+1}$ is considered a trusted prototype for two reasons: 1) In terms of magnitude, this is because the normalized malicious prototype do not affect the magnitude of trusted prototype, only the direction of trusted prototype; 2) In terms of direction, the average of all prototypes remains a plausible direction due to the small proportion of malicious clients. Subsequently, Aggregator obtains the plaintext $\|\textit{{C}}^{\prime k}_{t+1}\|$ for subsequent cosine similarity computation. To achieve this, it computes the inner product $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}\cdot\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ and sends it to the Verifier for decryption, then receives the decrypted result. Then, cosine similarity, a widely used measure of the angle between two vectors, is employed to measure the credibility of local prototype. If the direction of local prototype is similar to that of trusted prototype, its credibility is higher. Formally, Aggregator computes the credibility of local prototype by the following formula,
@@ 6 Analysis
where $G(\lambda,\eta,E)=-\left(\eta-\frac{\eta^{2}L_{1}}{2}\right)\sum_{e=\frac{1}{2}}^{E}\left\|\nabla\mathcal{L}_{m,t}^{(e)}\right\|_{2}^{2}+\frac{E\eta^{2}L_{1}}{2}\sigma^{2}+2\lambda.$
@@ 7 Experiments
7.1 Experimental Settings
7.1.1 Datasets and Models
@@
Furthermore, we apply CNN as local model to both MNIST and FMNIST. ResNet18 is used for CIFAR10, where we initialize the local model with pre-trained parameters. These initial parameters have an accuracy of 27.5% on CIFAR10’s test set.
7.1.2 Settings of FL
We employ a cross-silo configuration in FL setup. Specifically, we set up 20 clients, each of which uploads local prototypes at each communication round. The number of communication rounds is set to 100, 150, and 150 for MNIST, FMNIST, and CIFAR10, respectively. We configure the local learning rate to $\eta=0.01$, the importance weight to $\lambda=1$, a batch size to 64, and the number of local iterations to 5 by default. These hyperparameters are consistent across all clients.
7.1.3 Non-IID Settings
To simulate the Non-IID data in cross-silo FL, we create class-space heterogeneity among clients, which is common in the cross-silo scenarios. Specifically, large organizations (e.g., hospitals, companies) possess different data classes, and their class distributions may differ significantly, or even be missing some classes altogether. When these organizations participate in federated training, the union of their data classes defines the entire FL classification task. This phenomenon leads to Non-IID data across organizations. The Dirichlet distribution assumes that each client’s data is sampled from all classes, meaning each client typically have a certain percentage of all classes. However, in cross-silo scenarios, this contradicts the assumptions of Dirichlet distribution.
To model the data distribution in cross-silo scenarios, we define Avg as the mean number of data classes per client, and Std as the standard deviation of these class counts. In our experiments, we fix Avg to be 3 , 4 or 5, and fix Std to be 1 or 2, aiming to create the class-space heterogeneity. Clients are randomly assigned classes, with partial class overlap among them. To visualize the different data distributions, we plot heat maps as shown in Fig. 2.
@@ 7.1.3 Non-IID Settings
Fig. 2: Heat maps for different data distributions under heterogeneity distribution.
7.1.4 Setting of Data Poisoning Attacks
@@
7.1.5 Evaluation Measure
Our goal is to improve the performance of cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks. Therefore, we evaluate the FL performance by testing the average accuracy of benign clients. For experiment comparison, we use FedAvg as the baseline, assuming that FedAvg has not suffered data poisoning attacks. In addition, we compare our PPFPL with robust schemes (i.e., Krum [44] and Foolsgold [45]) and privacy-preserving robust schemes (i.e., ShieldFL [18] and PBFL [16]) under the same experiment conditions.
7.2 Experimental Results
7.2.1 Visualization of Prototypes
We use t-SNE to visualize client-submitted prototypes in PPFPL, as shown in Fig. 3. Specifically, we display prototypes from clients under feature and label attacks with Att =20% on the CIFAR10 dataset. This figure shows the distribution of these prototypes in a single communication round. We observe that the prototypes of malicious clients deviate significantly from those of benign clients, regardless of the features attacks or labels attacks. In addition, we notice that the prototypes submitted by malicious clients are concentrated in a ring area. This is because when malicious clients perform local model training, the features or labels are disrupted, causing the directions of the generated prototypes to spread from the center point to the surrounding areas. To mitigate the severe impact of malicious clients, it is essential to employ a secure aggregation protocol that ensures Byzantine-robust aggregation results. The following is the security evaluation of PPFPL.
((a)) Feature attacks, Att = 20%, $\textit{Avg}=3$, $\textit{Std}=2$.
((b)) Feature attacks, Att = 20%, $\textit{Avg}=5$, $\textit{Std}=1$.
((c)) Label attacks, Att = 20%, $\textit{Avg}=3$, $\textit{Std}=2$.
((d)) Label attacks, Att = 20%, $\textit{Avg}=5$, $\textit{Std}=1$.
Fig. 3: The t-SNE visualization of prototypes submitted by benign and malicious clients: different colors indicate distinct classes, where $\cdot$” denotes prototypes from benign clients, and “$\times$” represents those from malicious clients.
7.2.2 Security Evaluation
@@
Fig. 4: Test average accuracy comparison between PPFPL and existing schemes in poisoned Non-IID data.
@@
Fig. 5: Loss comparison between PPFPL and existing schemes in poisoned Non-IID data.
To evaluate the security of our framework, we test PPFPL’s performance against feature attacks and label attacks on three datasets under Non-IID setting with $\textit{Avg}=3$ and $\textit{Std}=2$. Specifically, the proportion of malicious clients is set to Att = 20% and the detection threshold of PPFPL is configured as $\chi=0$. Furthermore, we compare PPFPL with existing schemes (i.e., FedAvg, ShieldFL, and PBFL), where the global model learned by FedAvg is not subject to data poisoning attacks.
The experiment results are shown in Fig. 4, where the black dashed line represents the highest accuracy of PPFPL without attacks. We observe that the accuracies of ShieldFL and PBFL (under attacks) are severely lower than FedAvg (no attack), especially for CIFAR10. This indicates that as the dataset becomes more complex, the negative impact of data poisoning attacks on model accuracy becomes more obvious. This is mainly because the complex dataset increases the difficulty of defense, and poisoned Non-IID data degrades the defense performance. In contrast, PPFPL (under attack) outperforms FedAvg (no attacks), with its accuracy approaching the highest accuracy of PPFPL without attacks when it nears convergence. This demonstrates that PPFPL guarantees the high accuracy of the learned model in the presence of poisoned Non-IID data, which is benefited that the prototype is not affected by tampered data distribution, while the secure aggregation protocol resists malicious prototypes submitted by clients. Thus, our PPFPL overcomes the difficulty confronted by these defense schemes, and improves the FL performance in poisoned Non-IID data.
Additionally, we evaluate the change of loss during training for PPFPL and the existing schemes, as shown in Fig. 5. We notice that the loss of PBFL fluctuates significantly during training process, while PPFPL is relatively smooth. This indicates that data poisoning attacks disrupt the convergence of PBFL, while the convergence of PPFPL is not affected by data poisoning attacks, thus ensuring the accuracy of the model. This is consistent with our Corollary 3. Specifically, PPFPL still satisfies convergence in the presence of data poisoning attacks as long as $\lambda$, $E$, and $\eta$ satisfy a specific relationship among them. Moreover, PPFPL’s loss at convergence is lower than other defense schemes under CIFAR10, which can satisfy the design goal of security. Notably, our experiments show that the performance impact of feature attacks or label attacks is similar for PPFPL and other schemes. This indicates that the impact on model performance is similar under poisoned Non-IID data, regardless of the type of data poisoning attacks.
@@
Fig. 6: Stability of PPFPL under CIFAR10 against data poisoning attacks.
7.2.3 Different Data Distributions for PPFPL
To further test the security of PPFPL in different data distributions, we evaluate the accuracy of PPFPL against feature attacks under different Non-IID conditions in MNIST, FMNIST and CIFAR10. In addition, we compare the performance of Krum, Foolsgold, ShieldFL, and PBFL. The specific experimental results are shown in Table 1. Notably, since test average accuracy has fluctuated after each communication round, we are select the average of five highest test average accuracies across the communication rounds. From Table 1, we can observe that the performance of PPFPL is less affected by changes in data distribution, while other schemes are very susceptible to data distribution. This is attributed to that the client-submitted prototype does not change due to the change in data distribution. In addition, we observe a slight degradation in the performance of PPFPL with higher proportion of attacks. This is because the tampered Non-IID data reduces the contribution of effective samples to the model, resulting in a slightly lower performance, which is a reasonable phenomenon. Therefore, our PPFPL improves performance in cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks, which is applicable to deployments in real-world scenarios.
Table 1: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with feature attacks.
@@ 7.2.3 Different Data Distributions for PPFPL
7.2.4 Secure Aggregation Protocol for PPFPL
To test the usefulness of aggregation protocol, we evaluate the performance of PPFPL with $\chi$ of -1, 0, 0.2, and 0.5, respectively, where $\chi=-1$ means that the protocol perform normalization verification and average aggregation, and the data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$). In addition, the $\boldsymbol{X}$ in Table 2 denotes that client-submitted prototypes are unnormalized and aggregated by simple averaging. From Table 2, we can observe from $\boldsymbol{X}$ that the performance of PPFPL suffers degradation to some extent compared to other settings, but the magnitude of that degradation is limited. This is because each client minimizes its local classification loss and aligns local prototype with global prototype during local model training. However, the poisoned global prototype can only affect part of local model training, and cannot affect the minimization of local classification loss. Thus, in PPFPL, the model accuracy of benign clients is less susceptible to the influence of malicious clients. Furthermore, we observe that the presence of detection threshold improves the performance of PPFPL, which suggests that the secure aggregation protocol computes Byzantine-robust aggregation results, which reduces the influence of malicious clients and proves the effectiveness of the secure aggregation protocol. The above observation raises the question: whether PPFPL can maintain high performance in high proportion of attacks?
Table 2: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with diffierent $\chi$ on feature attacks.
@@ 7.2.4 Secure Aggregation Protocol for PPFPL
Fig. 7: Test average accuracy of different $\lambda$ with different proportions Att with feature attacks.
7.2.5 High Proportion Attacks for PPFPL
To test the performance of PPFPL under high proportion attacks, we evaluate the performance of PPFPL under Att = 20%, 40%, 60% and 80%, respectively. In addition, we set $\lambda$ as 0.01, 0.1 and 1 respectively, and the data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$). The $\lambda$ denotes the importance weight of the auxiliary term in loss function, which can be considered as the degree of influence among clients. The experimental results are shown in Fig. 7. We surprisingly observe that PPFPL still has high performance when Att is 60% or even 80%, which indicates that the training of benign clients is not interfered by malicious clients. This is because the benign client does not rely solely on information distributed by two servers in federated prototype learning, but relies heavily on its local training data. Furthermore, we observe that when Att is 20% or 40%, the accuracy increases slightly with the growth of $\lambda$ in the three datasets. On the contrary, when Att is 60% or 80%, the accuracy tends to decrease with the increase of $\lambda$. This is because the larger $\lambda$ strengthens the collaboration among clients, but makes PPFPL more vulnerable to data poisoning attacks from malicious clients. Conversely, the smaller value of $\lambda$ weakens collaboration among clients, but increases resistance to data poisoning attacks. Therefore, PPFPL can appropriately adjust the size of $\lambda$ according to actual conditions.
7.2.6 Stability of PPFPL

### v4

@@ V-B2 Prototype Handling
V-C Secure Aggregation
To resist local prototypes submitted by malicious clients without compromising privacy, we design a secure aggregation protocol across two servers to filter malicious local prototypes and obtain Byzantine-robust global prototypes. The protocol consists of normalization verification and secure two-party computation.
V-C1 Normalization verification
Since malicious clients may bypass the normalization stage to amplify the impact of their local prototypes, the two servers need to verify that encrypted local prototypes are normalized. Specifically, Aggregator calculates the inner product $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ for each local prototype, and sends the results to Verifier, where $\cdot$ denotes the inner product. Then, the Verifier decrypts $\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}\cdot\llbracket\widetilde{\textit{{c}}}^{k}_{m,t}\rrbracket_{Pkv}$ using its secret key $Skv$, and checks whether the inner product $\|\widetilde{\textit{{c}}}^{k}_{m,t}\|^{2}$ equals 1. If the inner product of prototype from client $m$ is not equal to 1, which indicates that its local prototype does not normalized, client $m$ is removed from the set of clients $\mathcal{S}$. After validation, Verifier sends the remaining client set $\mathcal{S}$ to Aggregator.
V-C2 Secure Two-party Computation
@@ 19       Distribute global prototype $\llbracket\textit{{C}}^{k}_{t+1}\rrbracket_{Pkx}$ to clients;
where $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ denotes the trusted prototype for class $k$ in $(t+1)$-th communication round. $\textit{{C}}^{\prime k}_{t+1}$ is considered a trusted prototype for two reasons: 1) In terms of magnitude, this is because the normalized malicious prototype do not affect the magnitude of trusted prototype, only the direction of trusted prototype; 2) In terms of direction, the average of all prototypes remains a plausible direction due to the small proportion of malicious clients. Subsequently, Aggregator obtains the plaintext $\|\textit{{C}}^{\prime k}_{t+1}\|$ for subsequent cosine similarity computation. To achieve the plaintext, it computes the inner product $\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}\cdot\llbracket\textit{{C}}^{\prime k}_{t+1}\rrbracket_{Pkv}$ and sends it to the Verifier for decryption, then receives the decrypted result. Then, cosine similarity, a widely used measure of the angle between two vectors, is employed to measure the credibility of local prototype. If the direction of local prototype is similar to that of trusted prototype, its credibility is higher. Formally, Aggregator computes the credibility of local prototype by the following formula,
@@ VI Analysis
where
@@ VII Experiments
VII-A Experimental Settings
VII-A1 Datasets and Models
@@
Furthermore, we apply Convolution Neural Network (CNN) as local model to both MNIST and FMNIST. For CIFAR-10, we employ ResNet18 as the local model, initialized with pre-trained parameters. These initial parameters have an accuracy of 27.5% on CIFAR10’s test set.
VII-A2 Hyper-parameters Settings of FL
We employ a cross-silo configuration in FL setup. Specifically, we set up 20 clients, each of which uploads local prototypes at each communication round. The number of communication rounds is set to 100, 150, and 150 for MNIST, FMNIST, and CIFAR10, respectively. We configure the local learning rate to $\eta=0.01$, the importance weight to $\lambda=1$, a batch size to 64, and the number of local iterations to $E=5$. Except for special declared, the above default settings are applied. These hyper-parameters are consistent across all clients.
VII-A3 Non-IID Settings
To simulate the Non-IID data in cross-silo FL, we create class-space heterogeneity among clients, which is common in the cross-silo scenarios. Specifically, large organizations (e.g., hospitals, companies) possess different data classes, and their class distributions may differ significantly, or even be missing some classes altogether. When these organizations participate in federated training, the union of their data classes defines the entire FL classification task. This phenomenon leads to Non-IID data across organizations. The Dirichlet distribution assumes that each client’s data is sampled from all classes, meaning each client typically have a positive percentage of every class. However, in cross-silo scenarios, some clients may have no samples from certain classes, which contradicts the assumptions of Dirichlet distribution.
To model the data distribution in cross-silo scenarios, similar to previous works [22][53], we define Avg as the mean number of data classes per client, and Std as the standard deviation of these class counts. In our experiments, we fix Avg to be 3 , 4 or 5, and fix Std to be 1 or 2, aiming to create the class-space heterogeneity. Clients are randomly assigned classes, with partial class overlap among them. To visualize the different data distributions, we plot heat maps as shown in Fig. 2.
@@ VII-A3 Non-IID Settings
Figure 2: Heat maps for different data distributions under heterogeneity distribution.
VII-A4 Setting of Data Poisoning Attacks
@@
VII-A5 Evaluation Measure
Our goal is to improve the performance of cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks. Therefore, we evaluate the FL performance by testing the average accuracy of benign clients. For experiment comparison, we use FedAvg as the baseline, assuming that FedAvg has not suffered data poisoning attacks. In addition, we compare our PPFPL with robust schemes (i.e., Krum [35], Foolsgold [54], and FedDMC [36]) and privacy-preserving robust schemes (i.e., ShieldFL [18], PBFL [16], and AntidoteFL [20]) under the same experiment conditions.
VII-B Experimental Results
VII-B1 Visualization of Prototypes
We visualize the prototypes submitted by clients in PPFPL using t-SNE, as shown in Fig. 3. Specifically, we display the distribution of prototypes from clients under feature and label attacks with $\textit{Att}=20\%$ on the CIFAR10 dataset within a single communication round. We observe that the prototypes of malicious clients deviate significantly from those of benign clients, regardless of the features attacks or labels attacks. In addition, we notice that the prototypes submitted by malicious clients are concentrated in a ring area. This is because when malicious clients perform local model training, the features or labels are disrupted, causing the directions of the generated prototypes to spread from the center point to the surrounding areas. Most importantly, we observe a clear difference between the prototypes submitted by benign and malicious clients. This difference is prominent because it is independent of the data distribution. The finding validates the effectiveness of using prototypes to defend against data poisoning attacks. To filter out these malicious prototypes, it is essential to employ a secure aggregation protocol that ensures Byzantine-robust aggregation results. The following is the security evaluation of PPFPL.
((a)) Feature attack,$\textit{Avg}=3,\textit{Std}=2$.
((b)) Feature attack,$\textit{Avg}=5,\textit{Std}=1$.
((c)) Label attacks, $\textit{Avg}=3$, $\textit{Std}=2$.
((d)) Label attacks, $\textit{Avg}=5$, $\textit{Std}=1$.
Figure 3: The t-SNE visualization of prototypes submitted by benign and malicious clients: different colors indicate distinct classes, where $\cdot$” denotes prototypes from benign clients, and “$\times$” represents those from malicious clients.
VII-B2 Security Evaluation
@@
Figure 4: Test average accuracy comparison between PPFPL and existing schemes in poisoned Non-IID data.
@@
Figure 5: Loss comparison between PPFPL and existing schemes in poisoned Non-IID data.
To evaluate the security of our framework, we test PPFPL’s performance against feature attacks and label attacks on three datasets under Non-IID setting with $\textit{Avg}=3$ and $\textit{Std}=2$. Specifically, the proportion of malicious clients is set to Att = 20% and the detection threshold of PPFPL is configured as $\chi=0$. Furthermore, we compare PPFPL with existing schemes (i.e., FedAvg, ShieldFL, PBFL, FedDMC, and AntidoteFL), where the global model learned by FedAvg is not subject to data poisoning attacks.
The experiment results are shown in Fig. 4, where the black dashed line represents the highest accuracy of PPFPL without attacks. We observe that the accuracy of ShieldFL and PBFL (under attacks) is severely lower than FedAvg (no attack) in CIFAR10. This indicates that as the dataset becomes more complex, the negative impact of data poisoning attacks on model accuracy becomes more obvious. This is mainly because the complex dataset increases the difficulty of defense, and poisoned Non-IID data degrades the defense performance. Notably, FedDMC and AntidoteFL achieve higher performance compared to FedAvg because they adjust the aggregation weights of the model updates, which mitigates the bias caused by Non-IID data and improves the performance of federated learning. Moreover, PPFPL (under attack) outperforms FedAvg (no attacks), with its accuracy approaching the highest accuracy of PPFPL without attacks when it nears convergence. This demonstrates that PPFPL guarantees the high accuracy of the learned model in the presence of poisoned Non-IID data, which is benefited that the prototype is not affected by tampered data distribution, while the secure aggregation protocol resists malicious prototypes submitted by clients. Thus, our PPFPL overcomes the difficulty confronted by these defense schemes, and improves the FL performance in poisoned Non-IID data.
Additionally, we evaluate the change of loss during training for PPFPL and the compared schemes, as shown in Fig. 5. We notice that the loss of PBFL fluctuates significantly during training process, while PPFPL is relatively smooth. This indicates that data poisoning attacks disrupt the convergence of PBFL, while the convergence of PPFPL is not affected by data poisoning attacks, thus ensuring the accuracy of the model. This is consistent with our Corollary 3. Specifically, PPFPL still satisfies convergence in the presence of data poisoning attacks as long as $\lambda$, $E$, and $\eta$ satisfy a specific relationship among them. Moreover, PPFPL’s loss at convergence is lower than other defense schemes under CIFAR10, which can satisfy the design goal of security. Notably, our experiments show that the performance impact of feature attacks and label attacks is similar for PPFPL and other schemes. This indicates that the impact on model performance is similar under poisoned Non-IID data, regardless of the type of data poisoning attacks.
@@
Figure 6: Stability of PPFPL under CIFAR10 against data poisoning attacks.
VII-B3 Different Data Distributions for PPFPL
To further test the security of PPFPL in different data distributions, we evaluate the accuracy of PPFPL against feature attacks under different Non-IID conditions in MNIST, FMNIST and CIFAR10. In addition, we compare the performance of Krum, Foolsgold, ShieldFL, PBFL, FedDMC, and AntidoteFL. The specific experimental results are reported in Table I. Notably, since test average accuracy has fluctuated after each communication round, we select the average of five highest test average accuracies across the communication rounds. From Table I, we can observe that the performance of PPFPL is less affected by changes in data distribution, while other schemes are very susceptible to data distribution. This is attributed to that the client-submitted prototype does not change due to the change in data distribution. In addition, we observe a slight degradation in the performance of PPFPL with $Att=30\%$ attacks. This is because the tampered Non-IID data reduces the contribution of effective samples to the model, resulting in a slightly lower performance, which is a reasonable phenomenon. Therefore, our PPFPL improves performance in cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks, which is applicable to deployments in real-world scenarios.
TABLE I: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with feature attacks.
@@ VII-B3 Different Data Distributions for PPFPL
VII-B4 Secure Aggregation Protocol for PPFPL
To test the usefulness of aggregation protocol, we evaluate the performance of PPFPL with $\chi$ of -1, 0, 0.2, and 0.5, respectively, where $\chi=-1$ means that the protocol perform normalization verification and average aggregation, and the data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$). In addition, the $\boldsymbol{X}$ in Table II denotes that client-submitted prototypes are unnormalized and aggregated by simple averaging. As shown in Table II, we can observe from $\boldsymbol{X}$ that the performance of PPFPL suffers degradation to some extent compared to other settings. The performance degradation stems from the influence of malicious clients’ prototypes on the aggregation process, leading to a poisoned global prototype. However, the poisoned global prototype can only affect part of local model training, and cannot affect the minimization of local classification loss. Thus, in PPFPL, the model accuracy of benign clients is less susceptible to the influence of malicious clients. Furthermore, we observe that the presence of detection threshold improves the performance of PPFPL, which indicates that the secure aggregation protocol computes Byzantine-robust aggregation results, which reduces the influence of malicious clients and proves the effectiveness of the secure aggregation protocol. The above observation raises the question: whether PPFPL can maintain high performance in high proportion of attacks? We explore this question in the next experiment.
TABLE II: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with diffierent $\chi$ on feature attacks.
@@ VII-B4 Secure Aggregation Protocol for PPFPL
Figure 7: Test average accuracy of different $\lambda$ with different proportions Att with feature attacks.
VII-B5 High Proportion Attacks for PPFPL
To test the performance of PPFPL under high proportion attacks, we evaluate it with Att = 20%, 40%, 60% and 80%, respectively. In addition, we set $\lambda$ as 0.01, 0.1 and 1 respectively, and the data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$). The $\lambda$ denotes the importance weight of the auxiliary term in loss function, which can be considered as the degree of influence among clients. The experimental results are shown in Fig. 7. We surprisingly observe that PPFPL still has high performance when Att is 60% or even 80%, which indicates that the training of benign clients is not interfered by malicious clients. This is because the benign client does not rely solely on information distributed by two servers in federated prototype learning, but relies heavily on its local training data. Furthermore, we observe that when Att is 20% or 40%, the accuracy increases slightly with the growth of $\lambda$ in the three datasets. On the contrary, when Att is 60% or 80%, the accuracy tends to decrease with the increase of $\lambda$. This is because the larger $\lambda$ strengthens the collaboration among clients, but makes PPFPL more vulnerable to data poisoning attacks from malicious clients. Conversely, the smaller value of $\lambda$ weakens collaboration among clients, but increases resistance to data poisoning attacks. Therefore, PPFPL can appropriately adjust the size of $\lambda$ according to actual conditions.
VII-B6 Stability of PPFPL